OpenStudy (anonymous):

determine the number of permutation of 18 items taken 5 at a time.

OpenStudy (anonymous):

18 chooses 5 = 1028160?

OpenStudy (anonymous):

ok thanks :)

OpenStudy (anonymous):

np

OpenStudy (cruffo):

I though the phase "18 choose 5" usually refers to combinations not permutations. Combinations: \[18C5 = \frac{18!}{5!(18-5)!}\] Permutations: \[18P5 = \frac{18!}{(18-5)!}\]